# Learning Fairer Representations with FairVIC

### v2

@@ Appendix C Lambda Ablation Study Results
Tables 6 and 7 show the full results for each model when the weights on the FairVIC terms are adapted. Table 6 shows the effect of changing $\lambda_{\text{acc}}$ while keeping the FairVIC terms equal so that $\lambda_{\text{var, inv, cov}}=\frac{1-\lambda_{\text{acc}}}{3}$, where $10<\lambda_{\text{acc}}<$ 1, and Table 7 sets $\lambda_{\text{acc}}=0.1$, and suppresses one or two FairVIC terms to explore the effect of only utilising one or two term(s) at a time. For full discussion and analysis of the results of the lambda ablation study, see Section 6.3.
@@
Adult Income 0.50 $0.1\overline{6}$     0.8506 $\pm$ 0.0052 0.6564 $\pm$ 0.0110     0.0630 $\pm$ 0.0117 0.0473 $\pm$ 0.0175 -0.1561 $\pm$ 0.0161 0.3686 $\pm$ 0.0301
@@
German Credit 0.50 $0.1\overline{6}$     0.7175 $\pm$ 0.0211 0.8029 $\pm$ 0.0199     0.1073 $\pm$ 0.0549 0.0745 $\pm$ 0.0406 -0.0851 $\pm$ 0.0605 0.8866 $\pm$ 0.0839

### v3

@@ Appendix C Lambda Ablation Study Results
Tables 7 and 8 show the full results for each model when the weights on the FairVIC terms are adapted. Table 7 shows the effect of changing $\lambda_{\text{acc}}$ while keeping the FairVIC terms equal so that $\lambda_{\text{var, inv, cov}}=\frac{1-\lambda_{\text{acc}}}{3}$, where $0<\lambda_{\text{acc}}<1$, and Table 8 sets $\lambda_{\text{acc}}=0.1$, and suppresses one or two FairVIC terms to explore the effect of only utilising one or two term(s) at a time. For full discussion and analysis of the results of the lambda ablation study, see Section 6.4.
@@
Adult
Income
0.50 $0.1\overline{6}$ 0.8506 $\pm$ 0.0052 0.6564 $\pm$ 0.0110 0.0630 $\pm$ 0.0117 0.0473 $\pm$ 0.0175 -0.1561 $\pm$ 0.0161 0.3686 $\pm$ 0.0301
@@
German
Credit
0.50 $0.1\overline{6}$ 0.7175 $\pm$ 0.0211 0.8029 $\pm$ 0.0199 0.1073 $\pm$ 0.0549 0.0745 $\pm$ 0.0406 -0.0851 $\pm$ 0.0605 0.8866 $\pm$ 0.0839